ID |
Date |
Author |
Type |
Category |
Subject |
4812
|
Mon Jun 13 19:26:42 2011 |
Jamie, Joe | Configuration | CDS | SUS binary IO chassis 2 and 3 moved from 1X5 to 1X4 |
While preping 1X4 for installation of c1lsc, we removed some old VME crates that were no longer in use. This freed up lots of space in 1X4. We then moved the SUS binary IO chassis 2 and 3, which plug into the 1X4 cross-connect, from 1X5 into the newly freed space in 1X4. This makes the cable run from these modules to the cross connect much cleaner. |
4811
|
Mon Jun 13 18:40:08 2011 |
Jamie, Joe | Update | CDS | Snags in the repair of LSC CDS |
We've run into a problem with our attempts to get the LCS control back up and running.
As reported previously, the Dolphin fiber connection between c1lsc and c1sus appears to be dead. Since we have no replacement fiber, the solution was to move the c1lsc machine in to the 1X4 rack, which would allow us to use one of the many available short Dolphin cables, and then use a long fiber PCIe extension cable to connect c1lsc to it's IO chassis. However, the long PCIe extension cable we got from Downs does not appear to be working with our setup. We tested the cable with c1sus, and it does not seem to work.
We've run out of options today. Tomorrow we're going to head back to Downs to see if we can find a cable that at least works with the test-stand setup they have there. |
4810
|
Mon Jun 13 16:27:10 2011 |
Jenne | Update | Electronics | Seismometer Box Update/graphs |
Quote from elog 4807: |
The noise graphs relating total noise of the Seismometer circuit (GURALP stuff) to the LIGO seismic noise curve have been completed started.
|
What Larisa meant to post (I'm sure) is something more like this (sorry it's a little squished...I put too many words in the legend):
I've only included the 2 noise contributions from the LISO model that seem to dominate the sum noise. The plot gets a little crazy if you include all of the non-important sources.

So, what's the point??
First, the new box design doesn't have any crazy-special op-amps in it, so the noise of the new box is probably comparable to the old box. So, if that's true, the old box may not have been limiting the differential seismic noise. This definitely needs to be checked out. I'll make a quickie LISO model of the old Guralp breakout box, to see what its noise actually looks like, according to LISO. If it wasn't ever the breakout box that was limiting us, what the heck was it??
Second, the current box design seems to be better than the Guralp Spec sheet noise by ~a factor of 10. It would be nice if that number were more like a factor of 100. Or at least 30. So some work needs to be done to find a lower-noise op amp for the voltage buffer (the first op amp in the circuit).
Next steps:
Since Larisa is now starting her SURF project with Tara and Mingyuan, I'll look into improving the design of this box by a factor of 3 or 10.
Then I'll need to make a mock-up of it, and test it out.
If successful, then I'll draw it up in Altium and have it made. Recall that there should be 2 outputs per seismometer channel, one with high gain, one with low gain. Then 3 seismometer channels per seismometer (X, Y, Z), and perhaps multiple seismometer inputs per box. So lots and lots of stuff all in the same box. It's going to be pretty cool. |
4809
|
Mon Jun 13 15:33:55 2011 |
Jamie, Joe | Update | CDS | Dolphin fiber between 1Y3 and 1X4 appears to be dead |
The fiber that connects the Dolphin card in the c1lsc machine (in the 1Y3 rack) to the Dolphin switch in the 1X4 rack appears to have died spontaneously this morning. This was indicated by loss of Dolphin communication between c1lsc and c1sus.
We definitively tracked it down to the fiber by moving the c1lsc machine over to 1X4 and testing the connection with a short cable. This worked fine. Moving it back to using the fiber again failed.
Unfortunately, we have no replaced Dolphin fiber. As a work around, we are stealing a long computer->IO chassis cable from Downs and moving the c1lsc machine to 1X4.
This is will be a permanent reconfiguration. The original plan was to have the c1lsc machine also live in 1X4. The new setup will put the computer farther from the RF electronics, and more closely mimic the configuration at the sites, both of which are good things. |
4808
|
Mon Jun 13 12:34:21 2011 |
Jamie, Koji | Update | LSC | Updated LSC model installed |
After a couple of hickups, I was able to compile and install Koji's rework of the LSC model.
The main changes are that the model now use an RF_PD library part, and the channel names were tweaked to be more in line with what we expect.
I found a couple of small bugs in the model that were preventing it from compiling. Those were fixed and it compiled with no further problems.
There was also some rearrangement of signal inputs to the PD_DOF matrix. The matrix screen was updated to reflect the proper inputs. However, this also meant that the burt restore scripts for the IFO configurations were setting the wrong elements in the matrix. I fixed the settings for X and Y arm locking, and updated the burt snapshots using the burt/c1ifoconfigure/C1save{X,Y}arm scripts. NOTE: burt settings will need to be updated for the MICH, PRM, DRM, and FULL IFO configurations as well.
During the build/install process, Joe and I also found a bug in the feCodeGen that was causing the filter screens to be created with the wrong names. Joe sent out a patch that will hopefully be merged soon. Building the model with Joe's patch fixed the screen names, so the screens are currently named correctly. |
4807
|
Fri Jun 10 20:23:56 2011 |
Larisa Thorne | Update | Electronics | Seismometer Box Update/graphs |
(continuation of this)
The noise graphs relating total noise of the Seismometer circuit (GURALP stuff) to the LIGO seismic noise curve have been completed started.
I apparently harbor hate towards Matlab (you may have notice I do everything in Mathematica)....I will try to change my ways DX |
Attachment 1: SeisNoiseGraphs.jpg
|
|
4806
|
Fri Jun 10 18:49:40 2011 |
kiwamu | Update | IOO | Intensity Noise after the MC |
Last night the relative intensity noise (RIN) of the beam after MC was measured.
It looks like the RIN is dominated by the motion of the MC mirrors, possibly the angular motions because we don't have any angular stabilization servos.
Suresh will estimate the contribution from the MC mirrors' angular motion to the RIN in order to see if this plot makes sense.
(RIN)
The spectrum below 30 Hz seems to be dominated by the motion of the MC suspensions because it very resembles the spectra of those.
Above 30 Hz the spectrum becomes somehow flat, which I don't know why at the moment.
A rough estimation of the shot noise gave me a level of 10^{-9} RIN/sqrtHz, which is way below the measured spectrum.

(Setup)
All of the suspended mirrors were intentionally misaligned except for the MC mirrors and PRM to avoid interference from the other optics.
In this setup it allows us to measure the intensity noise of the laser which is transmitted from MC.
The beam transmitted from MC is reflected by PRM and finally enters into the REFL11 RFPD.
The DC signal from the RFPD was acquired at C1:LSC-REFL_DC_DQ as the laser intensity.
As well as the RIN measurement I took a spectrum when the beam is blocked by a mechanical shutter on the PSL table.
This data contains the dark noise from REFL11 and circuit noise from the whitening, AA board and ADC. It is drawn in black in the plot.
The cut off at 15 Hz is possibly due to the digital unwhitening (two poles at 15 Hz and two zeros at 150 Hz) filter to correct the analog whitening filter. |
4805
|
Fri Jun 10 14:50:37 2011 |
Koji | Update | LSC | New LSC code installed |
[Koji Jamie]
The new c1lsc code has been installed. The LSC screens have also been updated (except for ASS screen).
The major changes are:
1. Naming of the RFPD channels. Now the PD signals were named like:
REFL11_I_IN1, REFL11_I_IN2, REFL11_I_OUT ....
instead of REFL11I_blah
2. NREFL11, etc has been removed. We now have the official error signals
named like
REFL11_I_ERR
We can't use the name "REFL11_I" for the error signal as this name is
occupied by the name of the filter module. |
4804
|
Fri Jun 10 12:04:57 2011 |
Jenne | Update | RF System | Bad RF connections!! |
I am in the process of calibrating AS55's shot noise, and I noticed that the AS55 PD input to the demod board was only finger-tight. I then checked all of the other SMA connections in the set of RF PD demod boards, and found several more that were loose, including all of the REFL55 connections. This is no good!!!! RF connections need to be tightened! I went through and tightened all of the offending connections with my personal Snap-on SMA wrench. |
4803
|
Fri Jun 10 12:02:10 2011 |
rana | Update | CDS | Second trends only go back 12 days |
Quote: |
While answering a quick question by Kiwamu, I noticed we only had second trends going back to 99050000 GPS time, May 27th 2011.
Trends (I thought) were intended to be kept forever, and certainly longer than full data, which currently goes back several months.
Jamie will need to look into this.
|
Our concept is to keep second trends for 1-2 months and minute trends forever. The scheme that Alan had worked out many years ago had it such that we could look back to 1998 and that the minute trends would be backed up somehow.
If its not working, we need to get Alan's help to recover the previous configuration. |
4802
|
Thu Jun 9 20:10:38 2011 |
kiwamu | Update | SUS | ETMY whitening filter : all off |
I checked the state of the whitening filters for the ETMY shadow sensors.
Result : They've been OFF (i.e. flat response).
(measurement and setup)
I measured the transfer functions of the whitening board (D000210) by looking at the signal before and after the whitening stage.
The whitening board handles five signals; UL, UR, LR, LL and SD, and there are five single-pin lemo outputs for each signal on the front panel.
A good thing on those lemo monitors is that their signals are monitored before the whitening stages.
Rana suggested me to use these signals for the denominator of the transfer functions and consider the sensor signals as excitation signals.
So I plugged those signals into extra ADC channels via an AA-board and measured the transfer functions.
In the measurement the coherence above 4 Hz was quite small while the suspension was freely swinging.
Therefore I had to excite the ETMY suspension by putting random noise in a frequency band from 5 Hz to 35 Hz to obtain better coherence.
(results)
The response is flat over frequency range from ~ 0.2 Hz to ~40 Hz, see the plot below.
According to the spectrum of each signal the measurements above 10 Hz are just disturbed by the ADC noise.
If the whitening filters are ON, a pole and zero are expected to appear at 30 Hz and 3 Hz respectively according to the schematic, but no such features.

|
4801
|
Thu Jun 9 18:25:22 2011 |
kiwamu | HowTo | CDS | look back a channel which doesn' exist any more |
For some purposes I looked back the data of some channels which don't exist any more. Here I explain how to do it.
If this method is not listed on the wiki, I will put this instruction on a wiki page.
(How to)
(1) Edit an "ini" file which is not associated to the real-time control (e.g. IOP_SLOW.ini)
(2) In the file, write a channel name which you are interested in. The channel name should be bracketed like the other existing channels.
example: [C1:LSC-REFL11_I_OUT_DAQ]
(3) Define the data rate. If you want to look at the full data, write
datarate = 2048
just blow each channel name.
Or if you want to look at only the trends, don't write anything.
(4) Save the ini file and restart fb. If necessary hit "DAQ Reload" button on a C1:AAA_GDS_TP.adl screen to make the indicators green.
(5) Now you should be able to look at the data for example by dataviewer.
(6) After you finish the job, don't forget to clean up the sentences that you put in the ini file because it will always show up on the channel list on dtt and is just confusing.
Also don't forget to restart fb to reflect the change. |
4800
|
Thu Jun 9 16:18:03 2011 |
josephb | Update | CDS | Second trends only go back 12 days |
While answering a quick question by Kiwamu, I noticed we only had second trends going back to 99050000 GPS time, May 27th 2011.
Trends (I thought) were intended to be kept forever, and certainly longer than full data, which currently goes back several months.
Jamie will need to look into this. |
4799
|
Thu Jun 9 12:21:07 2011 |
kiwamu | Update | PSL | Re : PMC needs help |
Nulling the slow actuation offset fixed the issue. Now PMC is back to normal.

The reflected beam on the CCD was quite symmetric (it looked very TEM00 mode !) for some reasons, I somehow suspected the mode matching to PMC.
One possibility I thought of was the laser temperature because it could change the laser spatial mode.
So I looked at the slow actuation offset on the FSS screen and found it was at -4.0 which sounds somewhat big.
Then I zeroed the offset by the slider and relocked PMC.
Then the spatial pattern of the reflected beam became usual (i.e. junk light looking) and the transmitted light wet up to 0.83 which is normal.
Quote from #4798 |
The PMC is losing power.
|
|
4798
|
Thu Jun 9 10:38:30 2011 |
steve | Update | PSL | PMC needs help |
The PMC is losing power. |
Attachment 1: pmcnothp.jpg
|
|
4797
|
Wed Jun 8 23:17:25 2011 |
kiwamu | Update | LSC | DRMI output matrix diagonalization |
Approximately two weeks ago I diagonalized the LSC output matrix for the DRMI locking.
Since actuation on the position of BS changes not only MICH but also PRC and SRC, we needed to diagonalize the output matrix.
- What I did :
(1) The DRMI was locked. At this point PRC, MICH and SRC was controlled by PRM, BS and SRC actuators respectively.
(2) I injected excitation signal on C1:LSC-MICH_EXC by awg. The excitation was at about 200 Hz, which is above the UGF of all the LSC loops.
At this point the excitation only shakes the position of BS.
(3) I looked at spectra of REFL11_I, AS55_Q, AS55_I, that were used to sense PRC, MICH and SRC respectively.
At the beginning I was able to see the peak due to the excitation in those spectra. This means BS shakes the other DOFs (i.e. PRC and SRC) as well as MICH.
(4) To minimize the coupling from MICH to PRC (or SRC), I tuned a number on an element of the output matrix, which transfers the signal from MICH to PRM (or SRM).
This business was done by looking at the peak on REFL11_I (or AS55_I) and minimizing it. Since this technique was too naive the tuning was done only in second decimal place.

|
4796
|
Wed Jun 8 22:48:09 2011 |
rana | Update | Cameras | ITM camera lenses changed |
I focused these lenses so that we could get a clean image of the mirrors and the OSEMs.
Our goal is to have an image where the optic diameter almost fills the entire monitor. We want the focus to be adjusted for the YAG beam (which is almost the same as focusing for the OSEMs). This will NOT produce a nice image of the cage using visible light, but that is just fine.
When Justin Garofoli was here he found a nice lens combo that did the job, so if anyone can find his old email or elog, lets just go back to that.
For now, we do not need a camera/lens system to focus very tightly on the center of the optic. |
4795
|
Wed Jun 8 16:41:48 2011 |
valera | Update | ASS | X and Y arm dither alignment status |
The current status of the dither alignment system:
- Both Xarm and Yarm alignment are working. The scripts are: scripts/autoDither/alignX(Y). Each script sets up the respective arm, turns on the dither lines and servos for 66 sec, offloads the control signals to TM alignment biases and PZT sliders in case of Yarm, and to TM and BS alignment biases in case of Xarm, and finally turns off and clears the servo filters and turns off the dither lines.
- Jammie witnessed the final tests of both scripts - both X and Y arm power went up from 0.6-0.7 to close to 1 and the AS beam became symmetric. Also Jammie wanted me to leave the ETMY oplev in its current non-nominal but more stable state i.e. the oplev signals go to the ADC from the D010033 card not the D020432 one. The scripts can now run from the CONFIGURE medm screen.
- Both arms use signals derived from modulating ITM and ETM in pitch and yaw dofs and demodulating the arm power (TRX or TRY) and the cavity length signal (AS55I). The Yarm actuation has 8 dofs - pitch and yaw of the ITM, ETM, and two input beam PZTs so all the sensed dofs are controlled. The Xarm actuation has only 6 dofs - pitch and yaw of the ITM, ETM, and BS. The Xarm servo is set up to servo the beam position on the ETMX and the relative alignment of the cavity and the input beam. The ITMX spot position is unconstrained and provides the null test. The residual displacement on the ITMX is 0.2-0.3 mm in yaw and 0.9-1.0 mm in pitch. The I phases of the beam centering lockins, which are also the error points of corresponding DOF filters, are calibrated in mm by unbalancing the TM coils by known amount. The attached snap shot of the medm screen now has both X and Y arm calibrated beam spot positions and uncalibrated input beam indicators. The input beam angle and position signals can/should be calibrated by tapping the signals digitally and applying the proper matrix transformation - this will require the model change.
- Currently there is no lock loss catching in the model. We should add a trigger on arm power (or an equivalent mechanism) to turn off the inputs to prevent the spurious inputs. |
Attachment 1: BeamPositionIndicators.png
|
|
4794
|
Tue Jun 7 16:11:09 2011 |
steve | Update | Cameras | ITM camera lenses changed |
Computar 75-12.5 zooms were installed for closer look at the resonant spots. Their alignment and focus needs more loving adjustment.
Atm 1, ITMX ( it was 10-40 mm Tamron before )
Atm 2, ITMY ( it was 12mm wide angle showing the towers before ) |
Attachment 1: P1070865.JPG
|
|
Attachment 2: P1070860.JPG
|
|
4793
|
Tue Jun 7 11:39:27 2011 |
Jamie | Update | SUS | No binary output module in ETMY |
Quote: |
1) The FM1 files in the XXSEN modules should switch the analog shadow sensor whitening. I found today that, at least on ETMY and ETMX, they do nothing. This needs to be fixed before we can use the suspensions.
|
Joe discovered today that ETMY in fact has no binary output module at all, so there is actually no digital control of the whitening filters at ETMY.
We suspect that the ETMY binary output module was maybe harvested to put in the LSC rack, but we're not sure.
We found a spare binary output adapter pcb, which I will try to assemble into a module to install in ETMY.
This does not explain what's going on with ETMX, though. ETMX has a binary output module, that appears to be properly hooked up. I'll try to debug what's going on there as well.
In the mean time, I've removed the ETMX binary output module to use as a reference for putting together another identical module for ETMY. |
4792
|
Mon Jun 6 23:56:16 2011 |
rana, valera | Configuration | SUS | ETMX/ETMY OSEM whitening |
We measured the OSEM PD whitening transfer function of the ETMX OSEM UL whitening stage (D000210) by comparing the input signal to the whitening amplifier (single pin LEMO monitor) to the output signal - both were piped into the DAQ. The transfer function was close to constant 0 dB/180 deg independent of the whitening switch selection (FM1 filter engaged/disengaged) up to ~20 Hz where we run out of coherence. All other ETMX and ETMY spectra at the input of the digital whitening compensation don't change when the whitening is switched on/off so by induction we conclude that all the ETMX/ETMY OSEM PD hardware whitening filters are not on. |
4791
|
Mon Jun 6 22:41:22 2011 |
rana | Update | SUS | Switching problem in SUS models |
Some weeks ago, Joe, Jamie, and I reworked the ETMY controls.
Today we found that the model rebuilds and BURT restores have conspired to put the SUS damping into a bad state.
1) The FM1 files in the XXSEN modules should switch the analog shadow sensor whitening. I found today that, at least on ETMY and ETMX, they do nothing. This needs to be fixed before we can use the suspensions.
2) I found all of the 3:30 and cts2um buttons OFF AGAIN. There's something certainly wrong with the way the models are being built or BURTed. All of our suspension tuning work is being lost as a consequence. We (Joe and Jamie) need to learn to use CONLOG and check that the system is not in a nonsense state after rebuilds. Just because the monitors have lights and the MEDM values are fluctuating doesn't mean that "ITS WORKING". As a rule, when someone says "it seems to work", that basically means that they have no idea if anything is working.
3) We need a way to test that the CDS system is working... |
Attachment 1: a.pdf
|
|
4790
|
Mon Jun 6 18:29:01 2011 |
Jamie, Joe | Update | CDS | COMPLETE FRONT-END REBUILD (WITH PROBLEMS (fixed)) |
Today Joe and I undertook a FULL rebuild of all front end systems with the head of the 2.1 branch of the RCG. Here is the full report of what we did:
- checked out advLigoRTS/branches/branch-2.1, r2457 into core/branches/branch-2.1
- linked core/release to branches/branch-2.1
- linked in models to core/release/src/epics/simLink using Joe's new script (userapps/release/cds/c1/scripts/link_userapps)
- remove unused/non-up-to-date models:
c1dafi.md c1lsp.mdl c1gpv.mdl c1sup_vertex_plant_shmem.mdl
- modified core/release/Makefile so that it can find models:
--- Makefile (revision 2451) +++ Makefile (working copy) @@ -346,7 +346,7 @@ #MDL_MODELS = x1cdst1 x1isiham x1isiitmx x1iss x1lsc x1omc1 x1psl x1susetmx x1susetmy x1susitmx x1susitmy x1susquad1 x1susquad2 x1susquad3 x1susquad4 x1x12 x1x13 x1x14 x1x15 x1x16 x1x20 x1x21 x1x22 x1x23 #MDL_MODELS = $(wildcard src/epics/simLink/l1*.mdl) -MDL_MODELS = $(shell cd src/epics/simLink; ls m1*.mdl | sed 's/.mdl//') +MDL_MODELS = $(shell cd src/epics/simLink; ls c1*.mdl | sed 's/.mdl//') World: $(MDL_MODELS) showWorld:
- removed channel files for models that we know will be renumbered
- For this rebuild, we are also building modified sus models, that are now using libraries, so the channel numbering is changing.
- make World
- this makes all the models
- make installWorld
- this installs all the models
- Run activateDQ.py script to activate all the relevant channels
- this script was modified to handle the new "_DQ" channels
- make/install new awgtpman:
cd src/gds make cp awgtpman /opt/rtcds/caltech/c1/target/gds/bin
- turn off all watchdogs
- test restart one front end: c1iscex
BIG PROBLEM
The c1iscex models (c1x01 and c1scx) did not come back up. c1x01 was running long on every cycle, until the model crashed and brought down the computer. After many hours, and with Alex's help, we managed to track down the issue to a patch from Rolf at r2361. The code included in that patch should have been wrapped in an "#ifndef RFM_DIRECT_READ". This was fixed and committed to branches/branch-2.1 at r2460 and to trunk at r2461.
- update to core/branches/branch-2.1 to r2460
- make World && make installWorld with the new fixed code
- restarted all computers
- restart frame builder
- burt restored to 8am this morning
- turned on all watchdogs
Everything is now green, and things seem to be working. Mode cleaner is locked. X arm locked.
|
4789
|
Mon Jun 6 17:28:31 2011 |
steve | Update | Computers | still, no PEM channels |
Quote: |
AA filter box was removed and modified at 1Y7 today. The -5V power supply was current limited when I plugged it back in. It was removed for medical attention.
NO PEM channels available! because of this.
|
D68L8EX-850Hz filter chips were removed and bypassed-shorted as Rana's entry on March 17, 2006 in old elog.
This unit is still has a short somewhere. |
Attachment 1: P1070846.JPG
|
|
4788
|
Mon Jun 6 17:22:09 2011 |
valera | Configuration | LSC | Clipping in the X arm 1064 um trans path |
I changed optics in the ETMX transmon path to remove clipping (which made a false QPD signal).
During the weekend I found that there was an offset in X arm c1ass pitch servo, which derives the signal by demodulating the arm cavity power, coming from the beam clipping in the transmon path.
The clipping was on the pair of the 1" mirrors that steer the beam after the 2" lens (see attached picture). The beam is about 5-6 mm in diameter at this distance from the lens and was not well centered.
I moved the steering mirrors downstream by about 8" where the beam is about 2-3 mm (the attached picture shows the mirrors in the new location). The Y arm layout is different from X arm and I didn't find any obvious clipping in transmon path.
The max X arm buildup went up from 1.3 to 1.5. I changed the TRX gain from -0.003 to -0.002 to obtain the normalized X arm power of 1 in this state. The MC refl DC is 1.6 out of 4.9 V and the Y arm buildup is ~0.9 so the TRX(Y) gains will have to be adjusted once the MC visibility is maximized. |
Attachment 1: XarmTransMon.pdf
|
|
4787
|
Mon Jun 6 16:44:34 2011 |
Koji | Update | Electronics | POY11 tested |
The full characterization of POY11 is found in the PDF.
Resonance at 11.03MHz
Q of 7.6, transimpedance 1.98kOhm
shotnoise intercept current = 0.17mA (i.e. current noise of 7pA/rtHz)
Notch at 21.99MHz
Q of 56.2, transimpedance 35.51 Ohm
Notch at 55.20MHz
Q of 48.5, transimpedance 37.5 Ohm
|
Attachment 1: POY11_test.pdf
|
|
4786
|
Mon Jun 6 02:09:39 2011 |
rana | Update | Electronics | POY11 Rework Nearly Complete |
I've finished tuning POY11 and it is now sitting on top of the analyzer waiting for Koji to test its noise.
Notes:
- R2 has been switched from a 50 -> 110 Ohm R and a (29 Ohm | 47 pF) in parallel to it. This makes the gain of the MAX4107 be ~20 above 100 MHz and ~5 around 11 MHz. The high gain at high frequency makes it stable (squashes the 200 MHz oscillation) and the low gain at low frequency is to prevent saturation (the raw 11 MHz transimpedance is too high).
- There are two notches: 22.12 MHz and 55.3 MHz.
- The TEST IN input has been disabled since it wasn't useful.
- 107 Ohms inserted into between the U11 output and the TSENSE feedthrough. This is to prevent oscillations when driving the long Dsub cable. This needs to be done on all Gold Box RFPDs.
- R4 -> 50 Ohms.
- Had trouble tuning this to get the resonance at 11.06 MHz. This turned out to be the parallel inductance coming from L3 (previously 1.45 uH) whereas we needed a total inductance of ~1.6 uH. So I changed L3 to 33 uH to get it to be negligible compared to 1.7 uH at 11 MHz. This needs to be considered for all the 11 MHz diodes.
|
Attachment 1: 777.png
|
|
4785
|
Sat Jun 4 15:26:04 2011 |
Larisa Thorne | Update | Electronics | Seismometer Box Update |
(continuation of this)
Here are the transfer function and noise plots of the seismometer box, using the op amps that are actually indicated on the original plan (THS4131, AD826). I added them to the LISO op amp library (can be found in /cvs/cds/caltech/apps/linux64/liso/filter/opamp.lib)
Next step is to compare the noise graph below to the seismic noise curve of the interferometer to verify that the seismometer box configuration won't affect the curve... |
Attachment 1: SeisBoxLISO_transfer.pdf
|
|
Attachment 2: SeisBoxLISO_noise.pdf
|
|
4784
|
Fri Jun 3 18:12:03 2011 |
steve | Update | Computers | no PEM channels |
AA filter box was removed and modified at 1Y7 today. The -5V power supply was current limited when I plugged it back in. It was removed for medical attention.
NO PEM channels available! because of this.
|
4783
|
Fri Jun 3 14:27:32 2011 |
Koji | Update | Electronics | POP55 test results |
The full characterization of POP55 is found in the PDF.
Resonance at 54.49MHz
Q of 2.5, transimpedance 241Ohm
shotnoise intercept current = 4.2mA (i.e. current noise of 37pA/rtHz)
Notch at 11.23MHz
Q of 2.4, transimpedance 6.2 Ohm
Notch at 110.80MHz
Q of 53.8, transimpedance 13.03 Ohm
|
Attachment 1: POP55_test.pdf
|
|
4782
|
Thu Jun 2 23:10:10 2011 |
Koji | Update | Electronics | REFL11 test results (Re: REFL11 fixed) |
The full characterization of REFL11 is found in the PDF.
Resonance at 11.062MHz
Q of 15.5, transimpedance 4.1kOhm
shotnoise intercept current = 0.12mA (i.e. current noise of 6pA/rtHz)
Notch at 22.181MHz
Q of 28.0, transimpedance 23 Ohm
Notch at 55.589MHz
Q of 38.3, transimpedance 56 Ohm
|
Attachment 1: REFL11_test.pdf
|
|
4781
|
Thu Jun 2 16:31:41 2011 |
Jamie | Update | CDS | aquired SUS channel name suffixes changed from _DAQ to _DQ |
CDS changed the suffix for all aquired channel names from _DAQ to _DQ. When we rebuilt the sus models, described in the previous log, the channel names were changed and consequently the channel files were completely rewritten.
To fix the issue, the latest archived channel file was copied back into the chans directory, and the suffixes were changed, as so:
cd /opt/rtcds/caltech/c1/chans
cp archive/C1SUS_110602_155403.ini C1SUS.ini
sed -i 's/DAQ/DQ/g' C1SUS.ini
We then restarted the models and the framebuilder. |
4780
|
Thu Jun 2 16:23:42 2011 |
Jamie | Update | SUS | SUS control models updated to use new sus_single_control library part |
A new library part was made for the single suspension controller (it was originally made from the c1scx controller), using the following procedure:
- Opened c1scx model (userapps/trunk/sus/c1/models/c1scx)
- Cut ETMX subsystem block out of SUS subsystem
- Pasted ETMX block into new empty library, and renamed it C1_SUS_SINGLE_CONTROL
- Tweaked names of inputs, and generally cleaned up internals (cosmetically)
- Saved library to: userapps/trunk/sus/c1/models/lib/sus_single_control.mdl
Once the new sus_single_control library part was made and the library was committed to the cds_user_apps repo, I replaced all sus controller subsystems with this new part, in:
- c1scx
- c1scy
- c1sus (x5 for each vertex mass)
All models were rebuild, installed, and tested, and everything seems to be working fine. |
4779
|
Thu Jun 2 10:19:37 2011 |
Alex Ivanov | Summary | DAQ | installed new daqd (frame builder) program on fb (target/fb/daqd) |
I hope that new daqd code will fix the problem with non-aligned at 16 seconds frame file GPS times.
I have compiled new daqd program under /opt/rtcds/caltech/c1/core/release/build/mx and installed it under
target/fb/daqd, then restarted daqd process on "fb" computer. It was installed with the ownership of user root
and I did chmod +s on it (set UID on execution bit). This was done in order to turn on some code to renice daqd process
to the value of -20 on the startup. Currently it runs as the lowest nice value (high priority).
controls@fb /opt/rtcds/caltech/c1/target/fb $ ls -alt daqd
-rwsr-sr-x 1 root controls 6592694 Jun 2 10:00 daqd
Backup daqd is here:
controls@fb /opt/rtcds/caltech/c1/target/fb $ ls -alt daqd.02jun11
-rwxr-xr-x 1 controls controls 6768158 Feb 21 11:30 daqd.02jun11
|
4778
|
Wed Jun 1 21:56:43 2011 |
Koji | Update | General | aLIGO BBPD / Green PD final result |
Here is the conclusive result for the circuit configuration for aLIGO BBPD and 40m Green PD.
- Use Mini-circuits MAR-6SM for the RF preamplifier. The 50Ohm input impedance is used for the RF transimpedance.
The maximum output is ~4dBm.
- Use Mini-circuits GALI-6 for the RF middle power amp. The gain is 12dB and the amplifier is linear up to +17dBm. i.e. This is still linear at the maximum output level of MAR-6SM.
- The total RF transimpedance is ~2k. The DC transimpedance is also 2k.
- The bandwidth is 80MHz with FFD100 and internal 25V bias. When S3399 is used, the bandwdith goes up to 180MHz
although the responsivity of FFD100 at 1064nm is better than S3399 by a factor of 1.5. At the 40m we will use S3399 for the green BB PD.
- By adding an LC network next to the PD, one of the unnecessary signal can be notched out.
As an example, 9MHz notch was placed for the FFD100 case.
- Noise level: ~10pA/rtHz as a floor noise level at around 30MHz. This corresponds to the equivalent dark current of 0.4mA.
Matt has finished the PCB layout. We will order small first batches, and stuff it for the test. Some of these will be the 40m green PD. |
Attachment 1: PD_response.pdf
|
|
Attachment 2: linearity_GALI_6.pdf
|
|
4777
|
Wed Jun 1 13:33:22 2011 |
koji, tara | Update | Electronics | TTFSS #7 |
We replaced GE81 by PZT2907A (PNP transistor) in TTFSS #7, it's working fine.
Last time I broke Q4 transistor, which is used in the low noise power module for TTFSS, (see the schematic) and could not find another PZT2907A, so GE81 was used temporarily. Now we changed it back to PZT2907A as designed. I tested it by checking the voltage outputs of the board. It works fine, all voltage outputs are correct. I labeled one of the slot on the blue cabinet tower and kept the rest of the transistors there.

Quote: |
I brought TTFSS set #7 to 40m and kept it in the electronic cabinet.
note that Q4 transistor has not been replaced back to PZT2907A yet. It's still GE82.
Q3 is now pzt3904, not PZT2222A.
|
|
4776
|
Wed Jun 1 11:31:50 2011 |
josephb | Update | CDS | MC1 LR digital reading close to zero, readback ~0.7 volts |
There appears to be a bad cable connection somewhere on the LR sensor path for the MC1 optic.
The channel C1:SUS-MC1_LRPDMon is reading back 0.664 volts, but the digital sensor channel, C1:SUS-MC1_LRSEN_INMON, is reading about -16. This should be closer to +1000 or so.
We've temporarily turned off the LRSEN filter module output while this is being looked into.
I briefly went out and checked the cables around the whitening and AA boards for the suspension sensors, but even after wiggling and making sure everything was plugged in solidly. There was one semi-loose connection, but it wasn't on the MC1 board, but I pushed it all the way in anyways. The monitor point on the AA board looks correct for the LR channels, although ITMX LR struck me as being very low at about -0.05 Volts.
According to data viewer, the MC1 LR sensor channel went bad roughly two weeks ago, around 00:40 on 5/18 UTC, or 17:40 on 5/17 PDT.
UPDATE:
It appears the AA board (or possibly the SCSI cable connected to it) is the problem in the chain. |
4775
|
Tue May 31 17:30:42 2011 |
steve | Update | SUS | ETMY damping restored |
ETMY's watch dogs were found tripped. They were restored. |
4774
|
Tue May 31 16:07:57 2011 |
Larisa Thorne | Configuration | Electronics | Seismometer Box Update |
(Continuation of this)
I plugged the circuit into the LISO program to generate the graphs below....the first graph is a plot of frequency (f, in Hz) versus gain (in dB), and frequency (f, Hz) versus phase (in degrees). Also included is the second graph, which is a noise plot of all circuit parts which contribute to the total noise of the circuit.
The only issue I had was that two of the op amps I'd picked (see third attachment for the original circuit diagram) for the circuit were not in LISO's op amp library. So I replaced THS4131 (from the voltage buffer part) and AD826 (from the ADC driver part) with AD797 and LT1037, respectively in order to generate the plots below....
There are notes calling the AD797 "ultra low noise, low distortion", whose data sheet can be found here: AD797
Notes also call LT1037 "low noise, high speed precision op amp", whose data sheet can be found here: LT1037
I've put these in temporarily only, as I don't know if they are appropriate choices for the job or even if we have them. Suggestions? |
Attachment 1: SeisBoxLISOplot1.pdf
|
|
Attachment 2: SeisBoxLISOplot2.pdf
|
|
Attachment 3: STS2diagram_original.pdf
|
|
4773
|
Tue May 31 15:45:37 2011 |
Jamie | Update | CDS | c1iscey IOchassis powered off for some reason. repowered. |
We found that both of the c1iscey models (c1x05 and c1scy) were unresponsive, and weren't coming back up even after reboot. We then found that the c1iscey IOchassis was actually powered off. Steve's accepts some sort of responsibility, since he was monkeying around down there for some reason. After powerup and reboot, everything is running again. |
4772
|
Tue May 31 14:29:00 2011 |
jzweizig | Update | CDS | frames |
There seems to be something strange going on with the 40m frame builder.
Specifically, there is a gap in the frames in /frames/full near the start of
each 100k second subdirectory. For example, frames for the following times are missing:
990200042-990200669
990300045-990300492
990400044-990400800
990500032-990500635
990600044-990600725
990700037-990700704
990800032-990800677
990900037-990900719
To summarize, after writing the first two frames in a data directory, the next ~10 minutes of frames are usually missing. To make matters worse (for
the nds2 frame finder, at least) the first frame after the gap (and all successive frames) start at an arbitrary time, usually not aligned to a 16-second boundary. Is there something about the change of directories that is causing the frame builder to crash? Or is the platform/cache disk too slow to complete the directory switch-over without loss of data?
|
4771
|
Tue May 31 11:34:13 2011 |
steve | Summary | SAFETY | safety glasses checked |
1064 nm transmison were measured of 40m safety glasses as shown . Their performance did not degrade. They are as good as their labels. |
Attachment 1: P1070823.JPG
|
|
4770
|
Tue May 31 11:26:29 2011 |
josephb | Update | CDS | CDS Maintenance |
1) Checked in the changes I had made to the c1mcp.mdl model just before leaving for Elba.
2) The c1x01 and c1scx kernel modules had stopped running due to an ADC timeout.
According to dmesg on c1iscex, they died at 3426838 seconds after starting (which corresponds to ~39 days). "uptime" indicates c1iscex was up for 46 days, 23 hours. So my guess is about 8 days ago (last Monday or Tuesday), they both died when the ADCs failed to respond quick enough for an unknown reason.
I used the kill scripts (in /opt/rtcds/caltech/c1/scripts/) to kill c1spx, c1scx, and c1x01. I then used the start scripts to start c1x01, then c1scx, and then finally c1spx. They all came up fine.
Status screen is now all green. I renabled damping on ETMX and it seems to be happy. A small kick of the optic shows the approriately damped response. |
4769
|
Mon May 30 23:14:27 2011 |
valera | Update | ASC | Y arm initial alignment |
I closed all 8 dither loops for the Y arm initial alignment: 2x2 centering servo (this worked before) and 2x2 input beam servo for both pitch and yaw.
So far it looks pretty good - the error points go to zero and the arm power goes up to 1.
The offloading to the alignment biases and the PZTs is not yet automated.
Today the PMC, MC, and Y arm were very cooperative and a pleasure to work with. |
4768
|
Fri May 27 17:52:53 2011 |
steve | Update | LSC | LSC rack cables strain relieved & labeled |
LSC rack 1Y2 cables are strain relieved and labeled. Spare and/or obsolete cables are laid out on the top of the beam tube and on the outside of the rack.
The POY 110 MHZ demodboard has a very touchy position in the VME crate. Watch out for it! It has to be fixed. |
4767
|
Thu May 26 17:10:21 2011 |
Jamie | Configuration | LSC | New digital lockin added to LSC model/screen |
Quote: |
A future improvement would be to modify the DCPD input matrix to be able to route those signals to the lockin as well. This is actually currently possible since the DCPD input matrix is just a subset of the full input matrix, but it's not available via medm yet.
|
I went ahead and added the lockin output to the DCPD input matrix.
|
4766
|
Wed May 25 20:12:55 2011 |
Jamie | Configuration | LSC | New digital lockin added to LSC model/screen |
Quote: |
This is OK....but, the input matrix should come from the same place as the regular input matrix: i.e. it should be just another row like CARM, DARM, etc. rather than have its own screen.
|
You're absolutely right. That was a brain-fart oversight. I fixed the model so that the input from the lockin comes from another output row in the RFPD input matrix. I then fixed the C1LSC medm screen accordingly:

This is obviously much simpler and more straight-forward.
A future improvement would be to modify the DCPD input matrix to be able to route those signals to the lockin as well. This is actually currently possible since the DCPD input matrix is just a subset of the full input matrix, but it's not available via medm yet. |
4765
|
Wed May 25 19:19:11 2011 |
Jamie | Configuration | CDS | !!!CHECK IN YOUR MODELS!!! |
!!!CHECK IN MODEL CHANGES!!!
Today I found three models that were modified, but not checked in to the SVN repository:
M sus/c1/models/lib/sus_single.mdl
M isc/c1/models/c1lsc.mdl
M isc/c1/models/c1mcp.mdl
I checked in the c1lsc model, since I think it was just the change that Kiwamu made in http://nodus.ligo.caltech.edu:8080/40m/4749. I left the others, since I have no idea what they are or who made them.
Please please please remember to commit your model changes to the SVN after you're done. This is particularly important for important models, such as c1lsc. If you don't check in your changes I can pretty much guarantee that at some point you will loose them.
|
4764
|
Wed May 25 19:03:59 2011 |
Jamie | Configuration | CDS | Update rtcds checkout of cds_user_apps with new top-level directory names. |
The top-level subsystem subdirectories in the cds_user_apps source repository were renamed today to be all lower case. This required checking out the new directory and updating all of the model links in /opt/rtcds/caltech/c1. Here is how I updated the cds_user_apps working tree:
cd /opt/rtcds/caltech/c1/userapps
mv trunk{,.bak}
svn update --depth=files trunk
svn update --depth=empty trunk/{cds,isi,isc,psl,sus}
I then fixed the links in the /opt/rtcds/caltech/c1/core/release/src/epics/simLink directory:
for link in $(find . -maxdepth 1 -type l); do ln -sf $(readlink $link | tr [:upper:] [:lower:]) ; done
A couple of things had to be cleaned up:
- /opt/rtcds/caltech/c1/userapps/trunk/cds/c1/models/c1uct.mdl was linked in, but that model doesn't seem to exist anymore, so I removed the link.
- a couple of things were linked from /opt/rtcds/caltech/c1/userapps/trunk instead of /opt/rtcds/caltech/c1/userapps/release, so I fixed those links.
- /opt/rtcds/caltech/c1/userapps/release/cds/test/models/llo/l1isctest.mdl was not checked out, so I checked it out and fixed the link (this model should really be named something different if it is of common use, or we plan on using it at the 40m).
|
4763
|
Mon May 23 18:16:42 2011 |
Koji | Configuration | LSC | POY55 installed |
The DC Transimpedance of POP55 was increased from 50 Ohm to 10010 Ohm. There is the offset of 46mV. This should be cancelled in the CDS.
Quote: |
Last Saturday the POY55 RFPD (see this entry) was installed on the ITMY optical bench for the trial of the DRMI locking.
Since the amount of the light coming into the diode is tiny, the DC monitor showed ~ 3 mV even when the PRC was locked to the carrier.
In order to amplify the tiny RF signal from the photo diode a ZHL amplifier was installed next to the RFPD. The RF amp is sitting on delrin posts for insulation from the table.
|
|